OpenStudy (anonymous):

Solve the system by the method of substitution. 2x-2y-z=-14 z=-8 2x+3y+3z=-6

OpenStudy (anonymous):

@vinnv226

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

Hmm, I would start by plugging in -8 for all the "z"s in the first and third equations, and then simplify. Since you know the value of a variable right off the bat, it makes sense to plug it in and make the other equations simpler

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

so when i pluged it in for the first equation i got 2x-16-8=-14

OpenStudy (anonymous):

is that wrong

OpenStudy (anonymous):

You plugged in -8 for the y and the z. You can only plug it in for "z" because thats what the second equation tells us. We don't know what y is yet

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

2x-2y=6

OpenStudy (anonymous):

Yep, exactly. So now plug in z=-8 to the third equation as well

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

2x+3y=18

OpenStudy (anonymous):

@vinnv226

OpenStudy (blurbendy):

for the first equation, if you substituted -8 in for z you should get: 2x - 2y - (-8) = -14 2x - 2y + 8 = -14 2x - 2y = -22

OpenStudy (anonymous):

o did i do the second right

OpenStudy (anonymous):

* second one

OpenStudy (blurbendy):

yup, looks good

OpenStudy (anonymous):

kk what do i do know

OpenStudy (blurbendy):

you can solve for another variable. for example, the second equation says: 2x+3y=18 so you could solve for x

OpenStudy (anonymous):

Yea, the first one was incorrect, and I didn't notice it. My mistake

OpenStudy (anonymous):

lol im confused

OpenStudy (anonymous):

no problem

OpenStudy (blurbendy):

you could solve for x for 2x+3y=18 you would start by subtracting 3y from both sides 2x = -3y + 18

OpenStudy (anonymous):

kk

OpenStudy (blurbendy):

you have get x by itself though, so what do you do with the 2 in front?

OpenStudy (anonymous):

okay i got 2x=15 now divide by 2 to get it by itself right

OpenStudy (blurbendy):

you can't add -3y and 18, the -3 is multiplied by y, so you can't add it to 18. but yes, you would divide by 2 x = -3/2y +18/2 x = -3/2y + 9

OpenStudy (anonymous):

lol sry

OpenStudy (anonymous):

so know what do i do with the -3/2y add it

OpenStudy (blurbendy):

no, we know x = -3/2y + 9 now, so we can substitute x in the first equation we found: 2x - 2y = -22 2(-3/2y + 9) -2y = -22

OpenStudy (anonymous):

-6y+18-2y=-22

OpenStudy (blurbendy):

not quite, 2 * -3/2 = -3 the two's cancel so you end up with: -3y + 18 - 2y = -22

OpenStudy (anonymous):

kk know do i add 2y to each side

OpenStudy (blurbendy):

you dont need to, you can just add -3y -2y first

OpenStudy (blurbendy):

what does that give you?

OpenStudy (anonymous):

wait i add them so they cancel but what about the other side will it be 18=-22+2y+3y

OpenStudy (anonymous):

add them toghehter u mean

OpenStudy (blurbendy):

no, i meant you can just evaluate -3y -2y on the same side first: -3y - 2y + 18 = -22 -5y + 18 = -22

OpenStudy (anonymous):

8

OpenStudy (blurbendy):

yes, now you know y = 8 we know y = 8 z = -8 so to find x we can just look at the first equation and solve for x: 2x - 2(8) -(-8) = -14

OpenStudy (anonymous):

x=2

OpenStudy (blurbendy):

2x -16 + 8= -14 2x -8 = -14 2x = -6 so what does x equal

OpenStudy (anonymous):

sry i got 20 its 22 added wrong answer -3

OpenStudy (blurbendy):

yes, x = -3

OpenStudy (blurbendy):

so, z = -8 y = 8 and x = -3

OpenStudy (blurbendy):

we are done!

OpenStudy (anonymous):

thank you

OpenStudy (blurbendy):

anytime